OpenStudy (anonymous):

The quotient of a number and four is twelve less than the number. What is the number? A. –4 B. 9.6 C. 8 D. 16

OpenStudy (anonymous):

so the set up is \[\frac{ x }{ 12 }=x-12\]

OpenStudy (cookie_monster):

D.

OpenStudy (anonymous):

sorry not 12, 4

OpenStudy (anonymous):

\[\frac{ x }{ 4 }=x-12\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

so the multiply both side by 4 to get x=4(x-12) distribute the 4 x=4x-48 get the x values on one sides -3x=-48 divide both sides by -3 x=16